approximately 10% of the glass bottles coming off a production line have serious flaws in the glass. If two bottles are randomly selected, find the mean and variance of the number of bottles that have serious flaws.
Answer by swincher4391(1107)

Let X be a binomial random variable with p=.10 and n = 2
E[X] (mean) = n*p = 2 * .10 = 0.20
Var[x] = n*p*q = 2*.10*.90 = 0.18